30cm of N2 is mixed with 40cm of H2 under suitable condition to produce NH3
1.Write the balanced chemical equation for the reaction
2. Assuming the reaction gives the maximum yield of NH3 calculate
The volume of NH3 produced
The excess of any
reactant gas
The total volume of gaseous mixture after reaction
1
Expert's answer
2021-10-10T10:24:46-0400

1. N2 + 3H2 → 2NH3

2. V(N2) = 30 mL

V(H2) = 40 mL

n(N2) "= \frac{30}{22.4 \times 10^{3}} = 1.34 \times 10^{-3} \;mol"

n(H2) "= \frac{40}{22.4 \times 10^{3}} = 1.78 \times 10^{-3} \;mol"

According to the reaction for each mole of N2 we need 3 moles of H2. So, H2 is a limitting reactant.

n(NH3) "= \frac{2}{3}n(H_2) = \frac{2}{3} \times 1.78 \times 10^{-3} = 1.19 \times 10^{-3} \;mol"

V(NH3) "= 1.19 \times 10^{-3} \times 22.4 \times 10^{3} = 26.66 \;mL"

The volume of NH3 produced is 26.66 mL.

"n(N_2)_{used} = \frac{1}{3}n(H_2) \\

= \frac{1}{3} \times 1.78 \times 10^{-3} \\

= 0.593 \times 10^{-3} \;mol"

"V(N_2)_{used} = 0.593 \times 10^{-3} \times 22.4 \times 10^3 = 13.3 \;mL"

ΔV(N2) = 30 -13.3=16.7 mL (excess of N2)

The total volume of gaseous mixture after reaction = 16.7 mL + 26.66 mL = 43.36 mL